OpenStudy (anonymous):

(Fill in box) A conveyor belt moves 450 ft in 3/8 h. Complete the unit rate. [ ] ft/h

OpenStudy (jhannybean):

Divide the feel by the hours.\[\frac{450 \ ft.}{\frac{3}{8} \ hr} = 450 \ ft \cdot \frac{8}{3} \ hr = 150\cancel{450} \ ft \cdot \frac{8}{\cancel{3}} \ hr = (150 \cdot 8) \frac{ft}{hr}\]

OpenStudy (anonymous):

Thank you!!!